Please can someone help me solve this problem:
Write the standard equation of the circle passing through (2,5) with the center at (-2,-3).

Answer by nerdybill(7384)

Write the standard equation of the circle passing through (2,5) with the center at (-2,-3).
.
Equation of a circle:
(x-h)^2 + (y-k)^2 = r^2
where
(h,k) is the center
r is the radius
.
The problem gives the center at (2,5)
Plugging that into our formula we have:
(x-2)^2 + (y-5)^2 = r^2
The only thing we need is the radius. Get this from the distance between (2,5) and (-2,-3):
d = sqrt[(x2-x1)^2 + (y2-y1)^2]
d = sqrt[(2+2)^2 + (5+3)^2]
d = sqrt[(4)^2 + (8)^2]
d = sqrt[16 + 64]
d = sqrt[100]
d = 10 (this is our radius)
.
so now we finally get:
(x-2)^2 + (y-5)^2 = 10^2